OpenStudy (anonymous):

A landscaper is designing a flower garden in the shape of a trapezoid. She wants the shorter base to be 3 yards greater then the height and the longer base to be 7 yards greater than the height. She wants the area to be 295 square yards. This is modeled by the equation: h^2+5h=296. Find the height that will give the desired area. Round to the nearest hundreth of a yard.

OpenStudy (jim766):

use the quadratic formula \[\frac{ -5+\sqrt{5^2-4(295)} }{ 2 }=14.86\] that is the height now just figure out the two bases. and use the area of trap formula A = 1/2h(b1 + b2) to check it